In a triangle if a = 8,b=10,c=12 and C=lambda A find the value of lambda.

Grade:11

1 Answers

Arun
5 years ago
use the law of cosines: 
c^2 = a^2 + b^2 - 2ab cos C 
cos C = [(144 - 64 - 100)/(- 160)] = 1/8 = 0.125 
arccos (0.125) = C 
Similarly if you find
Hence
Angle is c is twice of angle A
Hence lambda = 2